OpenStudy (anonymous):

The formula for degrees Celsius is: C = 5/9 (F - 32), where F stands for degrees Fahrenheit. Which of the following equations shows the formula correctly solved for F? 1.) F = 5/9 C + 32 2.) F = 9/5 C + 32 3.) F = 5/9 C - 32 4.) F = 9/5 C - 32

OpenStudy (anonymous):

You try to manipulate the formula. We have, C/5 = (F-32)/9 Therefore, 9C/5 = F-32 Or, F = 9C/5 + 32

OpenStudy (anonymous):

Which mathematical sentence represents the solution for a in the equation 5b = bc + 4a? 1.) a = 5b + bc/4 2.) a = 5b - bc 3.) a = 5b - bc/4 4.) a = 4a/5b - bc

OpenStudy (anonymous):

Thanks. :)

OpenStudy (anonymous):

@Spacemanifestation Think you could help me with this last one?

OpenStudy (anonymous):

You have, 5b = bc + 4a Keep in mind that we need to solve for a. Thus, 4a = 5b-bc Or, a = (5b -bc)/4

OpenStudy (anonymous):

Thanks! :)